The half-life of Yttrium-91 is 59 days. How much of a 90.3 g sample of Yttrium-91 will remain after 324 days?

(answer is in grams, no units required)
1
Expert's answer
2019-04-23T11:05:21-0400

Let’s use the formula for radioactive decay:


"N = N_0e^{- \lambda t},"

here, "N_0 = 90.3 g" is the initial amount of Yttrium-91, "N" is the amount of Yttrium-91 that will remain after 324 days, "\lambda = 0.693/T_{1/2}" is the decay rate, "T_{1\/2} = 59 days" is the half-life of Yttrium-91 and "t = 324 days" is the time elapsed.

Then, we get:


"N = N_0e^{- \lambda t} = 90.3 g \cdot e^{- \dfrac{0.693 \cdot 324 days}{59 days} } = 2.0 g."

Answer:

"N = 2.0 g."